Required information [The following information applies to the questions displayed below.] Fitness Fanatics is a regional chain of health clubs. The managers of the clubs, who have authority to make investments as needed, are evaluated based largely on return on investment (ROI). The company's Springfield Club reported the following results for the past year: Sales $ 1,400,000 Net operating income $ 70,000 Average operating assets $ 350,000 The following questions are to be considered independently. 2. Assume that the manager of the club is able to increase sales by $70,000 and that, as a result, net operating income increases by $18,200. Further assume that this is possible without any increase in average operating assets. What would be the club’s return on investment (ROI)?
Answer #1

New operating income = 70000+18200 = 88200

Return on investment = Operating income/Average operating assets = 88200/350000 = 25.20%